# Apparate: Rethinking Early Exits to Tame Latency-Throughput Tensions in ML Serving

### v1

@@ 5.3 Comparison with Existing EE Strategies
Apparate (ResNet50) 99.0-99.2% 40.9-88.6% -1.6-0.0%
BranchyNet 85.8-99.8% -11.0-88.3% -11.0%
BranchyNet+ 76.1-99.9% -11.0-88.3% -11.0%
BranchyNet-opt 99.0-99.7% -11.0-74.5% -11.0%
Apparate (BERT-base) 99.1-99.3% 9.1-24.1% 0.7-1.8%
DeeBERT 91.7-97.1% 13.2-36.1% -1.3-6.4%
DeeBERT+ 82.2-90.3% 31.7-36.1% 5.9-6.4%
DeeBERT-opt 99.0% 9.8-36.1% -1.4-6.4%
Table 2: Comparison with existing EE models. Results list ranges of accuracies or latency wins across all CV (top row) or NLP (bottom row) workloads. ‘+’ and ‘opt’ pertain to optimized tuning strategies described in §5.3.
We compare Apparate with two off-the-shelf EE models: BranchyNet [53] and DeeBERT [57]. BranchyNet extends ResNet models with ramps of the same style as Apparate, while DeeBERT extends BERT-base with deeper ramps (using the entire BERT pooler, as described in §3.1). For each, we follow their prescribed architectures, with ramps after every layer that are always active. We perform one-time tuning of thresholds as recommended by both works, and consider two variants: the default recommendation where all ramps must use the same threshold, and a more flexible version that removes this restriction (+). For both, threshold tuning is done optimally (via grid search), and is based on uniformly sampled data across the workload. For fair comparison, Apparate’s ramp budget is configured to support ramps at all layers (though it never does so).
Table 2 presents our results. The main takeaway is that existing EE approaches, even when favorably tuned, yield unacceptable drops in average accuracy up to 23.9% and 17.8% for CV and NLP. In contrast, Apparate consistently meets the imposed accuracy constraint (1% in this experiment) for both workloads. Further, even with such accuracy violations, tail latencies are 0.9-9.4% lower with Apparate than with these systems. The reason is again lack of adaptation: all ramps are always active despite their current efficacy which vary dramatically over time (§2.3), yielding undue overheads for large numbers of non-exiting inputs. In contrast, throughout these experiments, despite having a full ramp budget, Apparate maintained only 9.1-27.2% of all possible ramps.
@@
5.4 Microbenchmarks
To ease presentation, results in this section use representative CV and NLP models (ResNet50, GPT2-medium) running on a random corpus video and Amazon reviews, respectively. All reported trends hold for all considered workloads.
(a) ResNet50
(b) GPT2
Figure 18: Apparate’s wins for different accuracy constraints.
@@ 5.4 Microbenchmarks
Table 3: Apparate’s median latency wins vs. ramp budget.
Parameter sensitivity. Recall that Apparate ingests values for two key parameters: ramp aggression (i.e., a ramp budget) and accuracy constraint (i.e., acceptable accuracy loss). Fig. 18 and Tab. 3 studies the effect that these parameters have on Apparate’s latency wins. The findings are intuitive: Apparate’s latency savings over vanilla models decrease as ramp budgets shrink or accuracy constraints tighten. Both trends are a result of Apparate being granted less flexibility for adaptation, either via smaller acceptable threshold ranges (that meet accuracy targets), or active ramp capacity (and thus potential ramp configurations). Importantly, accuracy constraint has a larger impact on Apparate’s wins. The reason is that inter-ramp dependencies result in overlap in the set of inputs that can exit at any ramp when run in isolation; thus, wins from using more ramps eventually hits diminishing returns. Indeed, Apparate begins by using the full budget, only to quickly disable many ramps that have a net negative effect on serving latencies.
Ramp architectures. Although Apparate opts for using many lightweight ramps, it’s adaptation algorithms can support any ramp architecture. To illustrate this, we ran Apparate with DeeBERT’s more expensive ramps (described above). Overall, on the Amazon Reviews dataset, we find that these costlier ramps dampen Apparate’s latency savings by 4% since they constrain Apparate’s runtime adaptation in terms of feasible configurations, i.e., fewer active ramps at any time. Crucially, we note that accuracy constraints were still entirely met due to Apparate’s frequent threshold tuning.
@@
Table 4: Apparate on different serving platforms. Results show (median, p95) latency over vanilla models in ms.
@@
Impact of serving platform. Apparate runs atop existing serving platforms, responding to overall serving (and exiting patterns) rather than altering platform decisions, e.g., for queuing. Table 4 shows that, despite the discrepancies in platform scheduling strategies and knobs, Apparate’s performance wins are largely insensitive to the underlying platform when CV or NLP workloads are configured with the same SLO goal. For example, Apparate’s median latency savings for the Amazon workload and GPT-2 are within 2.9% when using Clockwork or TensorFlow-Serving.
@@
Profiling Apparate. Figure 11 in §3.2 analyzes the runtime and optimality of Apparate’s threshold tuning algorithm. Beyond that, Apparate includes two other overheads while running: ramp adjustment and coordination between its CPU controller and serving GPUs. Ramp adjustment rounds take an average of 0.5 ms. Coordination overheads are also low because of Apparate’s small ramp sizes (definitions and weights consume only $\sim$10KB) and profiling data (simply a top-predicted result with an error score, collectively consuming around $\sim$1KB). Thus, CPU-GPU coordination delays take an average of 0.5ms per communication, 0.4ms of which comes from fixed PCIe latencies in our setup.
@@
Importance of Apparate’s techniques. Apparate’s runtime adaptation considers frequent (accuracy-guided) threshold tuning, with periodic ramp adjustments. Table 2 highlights the importance of threshold tuning on average accuracies. Here, we evaluate the importance of ramp adjustment on Apparate’s latency improvements by comparing versions with and without this optimization. Overall, disabling ramp adjustment results in 20.8-33.4% lower median latency wins, though worst-case latency (and throughput) and accuracy constraints remain continually met.
6 Additional Related Work
A number of model-serving systems have been proposed [44, 17, 49, 22, 39, 7, 5, 4] where the focus is on serving large volumes of inference requests within a pre-defined SLO. Existing systems favor maximizing the system throughput while adhering to the latency constraints (2) by the use of intelligent placement [49, 22], batching [4] and routing [44]. To the best of our knowledge, no existing serving proposals focus on alleviating the latency-throughput tension.
The ML community has been actively working on early-exit networks, with several proposals focusing on the EE’s ramp architecture and exit strategy [57, 53, 28, 58, 36, 46, 64]. The architecture of the ramp depends on the domain, but it typically consists of one or more layers that provide information necessary to make an exit decision by emulating the original model. Replicating the last (few) layers is the common [57, 58], Apparate builds on this approach and prefers shallow ramps in its workflow (3). Once a ramp architecture is chosen, the exit strategy could be based on confidence of the labels [36] or entropy of the prediction [57]. More sophisticated approaches exist, e.g., instead of considering the ramps as fully independent, [64] uses counter-based exiting. Apparate’s focus is on leveraging EEs to resolve the latency-throughput tension in serving systems with a design that generalizes to a large class of EE architectures.
@@ 6 Additional Related Work
Optimizing model serving objectives based on workload characteristics has been discussed in recent works [16, 44, 63, 21, 49, 15, 62]. Inferline [16] optimizes cost in serving pipelines of models while adhering to strict latency constraints using intelligent provisioning and management. Shepherd [63] maximizes goodput and resource utilization in highly unpredictable workloads. Despite their impressive results, these works still optimize their metric of choice at the expense of latency and do not resolve the latency-throughput tension, which is the focus of our (complementary) work.
A recent line of work has focused on creating variants of an ML model to optimize serving performance. Some of these works look at execution graph level optimizations such as quantization and fusing to reduce inference latency [3, 1], while others replace the model with an equivalent one that meets the provided constraints. Solutions like Mystify [23] and INFaaS [44] generates and chooses model variants based on their intent and constraints (including performance). As shown in §5.2, Apparate’s wins persist even on compressed models, and is thus complementary to these works in that it can operate on their outputs. Finally, optimizing the execution of dynamic neural networks that alter NN execution (e.g., EEs, mixture of experts) was proposed in [18, 61]. These are low-level optimizations (e.g., at the GPU) which can benefit Apparate and improve its performance.
7 Conclusion
@@
We present Apparate, the first system that automatically injects and manages early exiting for ML inference. Key to Apparate’s ability to alleviate latency-throughput tensions in serving is its use of exiting only for fast results (not compute savings). This provides continual feedback on exits, and powers Apparate’s novel adaptation strategies for EE ramps and thresholds. Apparate lowers median latencies by 40.5-91.5% for CV and 10.0-24.2% for NLP workloads, while meeting accuracy constraints and preserving throughputs.
Appendix A Appendix
Input: $ramps$, list of ordered active ramp IDs
Input: $acc\_loss\_budget$, max accuracy loss tolerable
Input: $smallest\_step\_size$, smallest step size for incrementing thresholds
Output: $thresholds$, thresholds associated with each ramp
Output: $latency\_savings$, latency savings with searched thresholds
/* all thresholds start at 0, i.e. no EE */
$thresholds\leftarrow[0.0]*len(ramps)$ /* each ramp has its own step size */
1 $step\_sizes\leftarrow[smallest\_step\_size]*len(ramps)$ while True do
       $best\_ramp,overstepped\_ramps,latency\_savings\leftarrow pick\_ramp(ramps,thresholds,acc\_loss\_budget)$ /* find next ramp to update thresholds */
2       if $best\_ramp$ is valid then
             /* increment threshold of the selected ramp */
3             $thresholds[best\_ramp]\mathrel{+=}step\_sizes[best\_ramp]$
             /* double step_size */
4             $step\_sizes[best\_ramp]\mathrel{*}=2$
5      else
6             if $step\_sizes.all()\leq smallest\_step\_size$ then
7                   return $thresholds$, $latency\_savings$
8
      /* half step_size for overstepped ramps, double step_size for the rest */
9       for $ramp$ in overstepped ramps do
[…]
11
Algorithm 1 Threshold tuning algorithm
Model bs=1 Small Medium Large
ResNet18 9 1$\times$ 2$\times$ 4$\times$
ResNet50 22.7 1$\times$ 2$\times$ 4$\times$
ResNet101 52.8 1$\times$ 2$\times$ 4$\times$
VGG11 6.9 1$\times$ 2$\times$ 4$\times$
VGG13 8.2 1$\times$ 2$\times$ 4$\times$
VGG16 10.2 1$\times$ 2$\times$ 4$\times$
Distilbert-base 15.5 2$\times$ 4$\times$ 8$\times$
BERT-Base 29.4 2$\times$ 4$\times$ 8$\times$
BERT-Large 63.2 2$\times$ 4$\times$ 8$\times$
GPT2-medium 103 2$\times$ 4$\times$ 8$\times$
Table 5: Different SLOs used. All numbers are in ms, measured on the A6000.

### v2

@@ 4.4. Comparison with Existing EE Strategies
Apparate (ResNet50) 99.0–99.2% 46.6–88.6% -1.6–0.0%
BranchyNet 85.8–99.8% -11.0–88.3% -11.0%
BranchyNet+ 76.1–99.9% -11.0–88.3% -11.0%
BranchyNet-opt 99.0–99.7% -11.0–74.5% -11.0%
Apparate (BERT-base) 99.1–99.3% 13.7–14.7% 2.1–3.0%
DeeBERT 91.7–97.1% 13.2–36.1% -1.3–6.4%
DeeBERT+ 82.2–90.3% 31.7–36.1% 5.9–6.4%
DeeBERT-opt 99.0% 9.8–36.1% -1.4–6.4%
Table 2. Comparison with existing EE models. Results list accuracies and latency wins for all CV (top) or NLP (bottom) workloads. ‘+’ and ‘opt’ use the optimized tuning from §4.4.
Classification. We compare Apparate with two off-the-shelf EE models: BranchyNet (Teerapittayanon et al., 2017) and DeeBERT (Xin et al., 2020). BranchyNet extends ResNet models with ramps of the same style as Apparate, while DeeBERT extends BERT-base with deeper ramps (using the entire BERT pooler - §3.1). For each, we follow their prescribed architectures, with always-on ramps after every layer. We perform one-time tuning of thresholds as recommended by both works, and consider two variants: the default recommendation where all ramps use the same threshold, and a version that removes this restriction (+).
Table 2 presents our results. The main takeaway is that existing EE approaches, even when favorably tuned, yield unacceptable drops in average accuracy up to 23.9% and 17.8% for CV and NLP. In contrast, Apparate consistently meets the imposed accuracy constraint (1% in this experiment) for both workloads. Further, even with such accuracy violations, tail latencies are 0.9–9.4% lower with Apparate than these systems. The reason is lack of adaptation: all ramps are always active despite current efficacy, yielding undue overheads for non-exiting inputs. In contrast, throughout these experiments, despite having a full ramp budget (for fair comparison), Apparate maintained only 9.1–27.2% of all possible ramps.
@@
Generative. We compare Apparate with FREE (Bae et al., 2023), a state-of-the-art EE solution for T5-large that succeeded CALM (Schuster et al., 2022). FREE relies on a single, fixed ramp and fine-tunes the entire model to cater to that ramp. The ramp location and weights are selected once to maximize latency savings subject to a 1% accuracy constraint on a representative dataset (default to the first 3% of samples). Both FREE and Apparate use T5’s final decode head as the ramp architecture. Moreover, to manage tail latencies, Apparate (like FREE) uses a ramp budget of 1. The main difference is that Apparate dynamically adjusts the ramp position (and threshold) to maximize latency savings and preserve accuracy; FREE’s omission of runtime adaptation yields 5.5% accuracy losses in these experiments.
Despite FREE’s accuracy violations, Apparate brings median and 25th percentile TPT savings of 28.0–71.0% and 15.7–28.0% over FREE (Figure 18). 95th percentile latencies with Apparate are 1.8–2.4% lower than with FREE because Apparate regularly flushes a batch decoding once the ramp accumulates a pre-specified number of exited tokens.
4.5. Microbenchmarks
Results here use ResNet50 and GPT2-medium running on a random video and Amazon reviews for classification. Reported trends hold for all our workloads.
Figure 19. Apparate’s wins for different accuracy constraints.
@@ 4.5. Microbenchmarks
Table 3. Apparate’s median latency wins vs. ramp budget.
Parameter sensitivity. Recall that Apparate ingests values for two key parameters: ramp aggression/budget and accuracy constraint. Figure 19 and Table 3 studies the effect that these parameters have on Apparate’s latency wins. The findings are intuitive: Apparate’s latency savings over vanilla models decrease as ramp budgets shrink or accuracy constraints tighten. Both trends are a result of Apparate being granted less flexibility for adaptation. Importantly, accuracy constraint has a larger impact on Apparate’s wins. The reason is that inter-ramp dependencies result in overlap in the set of inputs that can exit at any ramp when run in isolation; thus, wins from using more ramps eventually hits diminishing returns.
Ramp architectures. Although Apparate opts for using many lightweight ramps, it’s adaptation algorithms can support any ramp architecture. To illustrate this, we ran Apparate with DeeBERT’s more expensive ramps (§4.4). Overall, we find that these costlier ramps dampen Apparate’s latency savings by 4% since they constrain Apparate’s runtime adaptation in terms of feasible configurations, i.e., fewer active ramps at any time. Crucially, we note that accuracy constraints were still entirely met due to Apparate’s frequent threshold tuning.
@@
Table 4. Apparate on different serving platforms. Results show (median, p95) latency over vanilla models in ms.
@@
Impact of serving platform. Apparate runs atop existing serving platforms, responding to serving and exiting patterns rather than altering platform decisions, e.g., queue management. Table 4 shows that, despite platform discrepancies, Apparate’s performance wins are largely insensitive to the underlying platform when CV or NLP workloads are configured with the same SLO goal. For example, median latency savings for the Amazon workload and GPT-2 are within 2.9% when using Clockwork or TF-Serving.
@@
Profiling Apparate. Figure 10 analyzes the run time and optimality of Apparate’s threshold tuning algorithm. Beyond that, Apparate includes two other overheads while running: ramp adjustment and communication between its CPU controller and GPUs for runtime monitoring. Ramp adjustment rounds take an average of 0.5 ms. Coordination overheads are also low (and non-blocking for serving - §3) because of Apparate’s small ramp sizes (definitions and weights consume $\sim$10KB) and profiling data (simply a top-predicted result with an error score, collectively $\sim$1KB). In total, CPU-GPU coordination delays take an average of 0.5ms per communication, 0.4ms of which comes from fixed PCIe latencies in our setup.
@@
Importance of Apparate’s techniques. Apparate’s runtime adaptation considers frequent (accuracy-guided) threshold tuning, with periodic ramp adjustments. Table 2 shows the importance of threshold tuning on average accuracies. We also evaluate the importance of ramp adjustment on Apparate’s latency wins by comparing versions with and without it. Overall, disabling ramp adjustment results in 20.8–33.4% lower median latency wins, though worst-case latency (and throughput) and accuracy constraints remain continually met.
5. Additional Related Work
Early exit networks. Existing proposals focus on exit ramp architecture and exit strategy (Xin et al., 2020; Teerapittayanon et al., 2017; Huang et al., 2018a; Xin et al., 2021; Liu et al., 2020; Schwartz et al., 2020; Zhou et al., 2020; Bae et al., 2023; Schuster et al., 2022) for specific models. Ramp architectures are domain-specific, but replicating the last (few) layers is the common practice (Xin et al., 2020, 2021); Apparate builds on this and prefers shallow ramps (§3). Existing exit strategies consider confidence of the labels (Liu et al., 2020), entropy of the prediction (Xin et al., 2020), or more sophisticated elements such as counters across ramps (Zhou et al., 2020). Apparate is agnostic to EE technique, and instead focuses on bringing EEs to arbitrary models and providing runtime management to maximize latency savings and accuracy preservation.
Model optimizations. Much work has focused on creating variants of ML models to optimize serving. Some employ compiler techniques to analyze (and optimize) execution at a graph or operator level for lower latency (ten, 2023; tvm, 2023). Other techniques compress models into versions with fewer layers or less weight specificity while adhering to accuracy objectives, e.g., quantization (Frantar et al., 2023; Kim et al., 2024; Lin et al., 2023), distillation (Huang et al., 2022; Sanh et al., 2019), and model pruning (Frantar and Alistarh, 2023; Ma et al., 2023). These efforts are complementary to Apparate, which would bring input-level compute adaptation to their outputs, i.e., on optimized or compressed models. Moreover, unlike many of these techniques, Apparate does not alter the original model’s weights and instead preserves its full predictive power for hard inputs and constant feedback.
@@ 5. Additional Related Work
Inference optimizations. Recent work optimizes model serving objectives based on workload characteristics (Crankshaw et al., 2020; Romero et al., 2021; Zhang et al., 2023b; Gujarati et al., 2017; Shen et al., 2019; Choi et al., 2021; Zhang et al., 2019). For instance, Inferline (Crankshaw et al., 2020) optimizes serving cost while adhering to strict latency constraints using intelligent provisioning and management. Shepherd (Zhang et al., 2023b) maximizes goodput and resource utilization by leveraging cross-workload predictability. Despite their impressive results, these works optimize their metric of choice at the expense of latency and do not resolve the latency-throughput tension, which is the focus of our (complementary) work. Mystify (Guo et al., 2021) and INFaaS (Romero et al., 2021) generate and choose model variants based on their intent and constraints (including performance). As noted above, Apparate operates on the output of such tools, bringing latency wins to compressed models (§4.2).
Dynamic neural networks. Other techniques boost inference efficiency by adapting model execution (Wang et al., 2018; Fedus et al., 2022; Huang et al., 2018b; Xu et al., 2018; de Vries et al., 2017) at different granularities, e.g., sample-wise, spatial-wise, and temporal-wise. As with exiting, their modulations to serving risk accuracy violations and foregone latency wins with dynamic workloads. We hope that Apparate can motivate and guide adaptation systems to navigate accuracy-latency-throughput tradeoffs for such techniques, but leave that to future work. Other low-level (e.g., GPU kernel level) techniques that optimize execution of dynamic neural networks (Cui et al., 2023; Zhang et al., 2023a) can benefit Apparate and improve its performance.
6. Conclusion
@@
We present Apparate, the first system that automatically injects and manages early exiting for ML inference. Key to Apparate’s ability to alleviate latency-throughput tensions in serving is its use of exiting only for fast results (not compute savings). This provides continual feedback on exits, and powers Apparate’s novel adaptation strategies for EE ramps and thresholds. Apparate lowers median latencies by 40.5–91.5% and 10.0–24.2% for diverse CV and NLP workloads, and reduces median time-per-token latencies by 22.6–77.9% for generative workloads, all while meeting accuracy constraints and preserving platform throughputs.
Acknowledgements. We thank Wyatt Lloyd, the SOSP reviewers, and our shepherd, Rong Chen, for their helpful and constructive comments. This work was supported by a Sloan Research Fellowship and NSF CNS grants 2147909, 2151630, 2140552, 2153449, and 2152313.